OpenStudy (anonymous):

im really bad with this slope intercept form :( write an equation in slope intercept form for the line that satisfies the following condition slope 7/18 and passes through (4,-20)

OpenStudy (mathstudent55):

The solpe-intercept form is y = mx + b where m = slope, and b is the y-coordinate of the y-intercept which is point (0, b)

OpenStudy (mathstudent55):

You were given that the slope is 7/18. That means you already know m = 7/18, and you can substitute that into the slope-intercept form: y = (7/18)x + b Now all you need to do is find b. Since you know the coordinates of a point, (4, -20), just substitute 4 into x and -20 into y in the equation above and solve for b: y = (7/18)x + b -20 = (7/18)*4 + b -20 = 28/18 + b -20 = 14/9 + b Multiply both sides by 9 to get rid of the denominator: -180 = 14 + 9b -194 = 9b 9b = -194 b = -194/9 Now that we know b, just substitute its value into the equation above: y = (7/18)x + b y = (7/18)x - 194/9